what are the x-interceps of y= 6x2 + 18x - 60
Answer by jim_thompson5910(35256)

Plug in y = 0 and solve for x.

y = 6x^2 + 18x - 60

y = 6(x^2 + 3x - 10)

0 = 6(x^2 + 3x - 10)

0/6 = x^2 + 3x - 10

0 = x^2 + 3x - 10

x^2 + 3x - 10 = 0

(x + 5)(x - 2) = 0

x + 5 = 0 or x - 2 = 0

x = -5 or x = 2

So the x-intercepts, as ordered pairs, are: (-5,0) and (2,0)